One of the ways to see what the predominant energy carrier produced is to see what was created.

So calculate the ATP, NADH, FADH2, and GTP that is created when your cell has consumed 3 glucose molecules and 1-10Carbon long fatty acid. Use these molecules until they have been fully consumed by the citric acid cycle.
Answer #1

For 1 glucose molecule oxidation-

Glycolysis- 2 ATP, 2 NADH

Pyruvate oxidation - 2 NADH

Krebs cycle-- 6 NADH, 2 FADH2, 2ATP/GTP

So for 3 glucose- total-6 ATP, 30 NADH, 6 FADH2 and 6 ATP/ GTP.

To break 10 carbon fatty acid 4 cycle of beta oxidation is required to produce 5 acetyl CoA and each cycle produces 1 NADH and 1 FADH2 so for 4 step 4 NADH and 4 FADH2.

But when 1 acetyl CoA enters into citric acid cycle it produces 3 NADH, 1 FADH2, and 1 ATP/ GTP.

SO for 5 acetyl CoA- 15 NADH, 5 FADH2 and 5 ATP/ GTP is produced.

So total NADH for fatty acid oxidation =15+4 = 19

FADH2. =4+5 = 9. and 5 ATP/ GTP

Now glucose+ fatty acid =6 ATP, 39 NADH, 15 FADH2, 11 ATP/ GTP.